3. Given the following information:

Quantity

Average Fixed Cost

Average Variable Cost

Average Total Cost

Marginal Cost

1

300

100

400

100

2

150

75

225

50

3

100

70

170

60

4

75

73

148

80

5

60

80

140

110

6

50

90

140

140

7

43

103

146

180

8

38

119

157

230

9

33

138

290

10

30

160

360

a) If market price is $140, then…

1. the firm will produce quantity ___________

2. Total revenue = _____________

3. Total cost = _______________

4. Profit for this firm = _______________

b) If market price is $230, then…

1. the firm will produce quantity ___________

2. Total revenue = _____________

3. Total cost = _______________

4. Profit for this firm = _______________

c) If market price is $80, then…

1. the firm will produce quantity ___________

2. Total revenue = _____________

3. Total cost = _______________

4. Profit for this firm = _______________
Answer #1

A) The firm will produce the quantity = 6

Total revenue = 140* 6 = 840

Total cost = 140*6 = 840

Profit = 840 - 840 = 0

B) the firm will produce the quantity = 8

Total revenue = 230*8 = 1840

Total cost = 157*8 = 1256

Profit = 584

C) the firm will produce the quantity = 4

Total revenue = 80*4 = 320

Total cost = 148*4 = 592

Profit = - 272